Project

General

Profile

Operations Readiness To Do List

Beam Line
Issue Responsible Party System Issue Description Data Noted Status Date Completed
MDR1-BL-01 C.MacWaters PPS A PPS module in the trench was faulty; this would have prevented the experiment from operating 2015-01-21 Wrongly reported, there was no problem 2015-03-16
MDR1-BL-02 C.MacWaters PPS The PPS mimic should be tested at a time when the ISIS PPS system is running and correct operation ensured 2015-01-21 This refers to MDR1 happening when ISIS server was OFF!. Mimic is tested all ok 2015-03-16
MDR1-BL-03 H.Nebrensky Target The target optical fibre system should be made more robust or appropriate operational constraints put in place to prevent interference with the target fibres 2015-01-21 Fibres to be fully enclosed eventually
Requirement that underfloor access banned when target active
October 2015
2015-03-02
MDR1-BL-04 H.Nebrensky Target The cause of the USB daemon crash on the target machine should be understood or appropriate operational measures put in place to ensure that it does not interfere with data taking 2015-01-21 Ensure this is covered in BLOC training 2015-03-02
MDR1-BL-05 P.Hodgson Target The cause of the target stepper motor driving against the limit switch should be understood and fixed 2015-01-21 Problem not reproduced 2015-02-06
MDR1-BL-06 H.Nebrensky BeamStop The beamstop operation should be checked in the light of a potential hydraulics leak 2015-01-21 Repair due on April maintenance day 2015-08-19
Detectors
Issue Responsible Party System Issue Description Data Noted Status Date Completed
MDR1-DET-01 P.Soler Luminosity Monitors The luminosity monitor should be connected to its power supply correctly 2015-01-21
MDR1-DET-02 P.Soler CKOV & GVa1 The replacement for SY127 power supply should be installed 2015-01-21
MDR1-DET-03 P.Soler ToF 0/1/2 & KL A replacement for SY527 power supply or appropriate spares should be acquired 2015-01-21
MDR1-DET-04 M.Uchida Tracker The waveguides and cabling on the tracker should be completed 2015-01-21 Waiting for completion of south PRY install before progressing
MDR1-DET-05 M.Uchida Tracker Sufficient spares for the tracker readout should be commissioned so that they are hot-swappable 2015-01-21 In progress
MDR1-DET-06 M.Uchida Tracker The tracker cryogenics temperature fine control should be repaired 2015-01-21
MDR1-DET-07 M.Uchida Tracker The tracker cryostats should be grounded properly in order to prevent excess electronics noise in the tracker readout 2015-01-21 In progress, final solution when cryostats back in place
MDR1-DET-08 M.Uchida Tracker The tracker should be made light tight 2015-01-21
MDR1-DET-09 M.Uchida Tracker Sufficient spares should be acquired for the tracker power supplies 2015-01-21 1 new spare, another on order 2015-03-16
MDR1-DET-10 M.Uchida Tracker The tracker internal and external LED systems should be made operational 2015-01-21
MDR1-DET-11 F.Drielsma EMR Investigate the noisy EMR Front End Boards and make some spares 2015-01-21 No overflow, spares in production 2015-06-23
MDR1-DET-12 F.Drielsma EMR Replace the faulty EMR SAPMT 2015-01-21 SAPMT readout fixed 2015-04-30
MDR1-DET-13 F.Drielsma EMR Implement and test the EMR LED Low Voltage PSU 2015-01-21 New LVPSU installed 2015-04-01
MDR1-DET-14 M.Uchida Tracker Tracker Cryo 3 power supply should be repaired 2015-01-21
MDR1-DET-15 M.Uchida Tracker Tracker Cryo 2 is leaky and should be repaired 2015-01-21
MDR1-DET-16 DL Team General Develop a global grounding scheme for the MICE hall 2015-01-21
Data Acquisition
Issue Responsible Party System Issue Description Data Noted Status Date Completed
MDR1-DAQ-02 Y.Karadzhov DATE A new unpacking should be released to online monitoring and online reconstruction developers 2015-01-21 Completed 2015-02-25
MDR1-DAQ-03 Y.Karadzhov DATE The crashes in the InfoBrowser should be understood and fixed. This is the principal tool for DAQ debugging and is needed for both shifters and experts 2015-01-21
MDR1-DAQ-04 M.Uchida Tracker DATE-tracker integration code should be cleaned up and validated 2015-01-21
MDR2-DAQ-01 Y.Karadzhov DAQ fADC V1724 geo13 continuously gets errors, investigate and if needed, replace the board with one of the spares 2015-03-05 Completed 2015-3-14
MDR2-DAQ-02 Y.Karadzhov DAQ Readout log files to be copied at the end of the run in the place where the binary data files are. (miceraid5:/data/mice/) 2015-03-05 Completed 2015-3-14
DAQ-01 Y.Karadzhov DAQ-EPICS Integrate the scaler with the new Trigger System. Ensure that the information required by the controls software is made available 2015-01-21 In progress
}
Controls & Monitoring
Issue Responsible Party System Issue Description Data Noted Status Date Completed
MDR1-CAM-01 P.Hanlet Alarms The alarm handler should be displayed on the alarm handler display and should be hooked up to the control room speakers 2015-01-21 Complete 2015-03-02
MDR1-CAM-03 P.Hanlet EMR Remote operation of the EMR should be established 2015-01-21 In progress
MDR1-CAM-04 M.Uchida Tracker Remote operation of the tracker should be integrated with the rest of the controls software 2015-01-21 In progress
MDR1-CAM-05 P.Hanlet Diffuser The "diffuser set" indicator should be fixed in the diffuser controls GUI 2015-01-21 In progress
MDR1-CAM-06 S.Boyd ISIS ISIS Beam Loss monitoring was not available. It should be determined whether this is part of the ISIS plan, and if so how MICE can monitor ISIS beam loss in the future 2015-01-21 Complete 2015-03-01
MDR1-CAM-07 P.Hanlet Diffuser/pAbsorber Improve mask control of diffuser and proton absorber 2015-01-21 Complete 2015-02-22
MDR1-CAM-08 H.Nebrensky BeamLine Provide state machine spreadsheet for beamline specifying beamline alarm limits, etc. 2015-01-21 Complete 2015-04-16
MDR1-CAM-09 P.Hanlet Alarms Implement and test beamline alarm handler and archiver settings; fix transition to offline state; same as CAM-08 2015-01-21 Complete 2015-04-16
MDR1-CAM-10 P.Hanlet RunControl Clean up unnecessary run control menus 2015-01-21 Complete 2015-02-17
MDR1-CAM-11 P.Hanlet RunControl Add buttons to Arm/Start/Pause/Stop DATE 2015-01-21 Near completion; still needs lock files
MDR1-CAM-12 P.Hanlet RunControl Consider adding Online Monitoring and Online Reconstruction notifications 2015-01-21 In progress
MDR1-CAM-13 P.Hanlet RunControl Confirm configuration write to CDB after first trigger, Begin/End run comments and tag run 2015-01-21 Requires beam
MDR1-CAM-14 P.Hanlet Target Add target optical readouts to Alarm Handler 2015-01-21 Awaiting Matt's fixing PV updating
MDR1-CAM-15 P.Hanlet DS Implement simplified shifter Decay Solenoid control panel 2015-01-21 Complete 2015-02-27
Computing
Issue Responsible Party System Issue Description Data Noted Status Date Completed
MDR1-CMP-01 P.Smith Web cams Need Konqueror working on all opi/thin PCs 2015-01-21 #1644
MDR1-CMP-02 R.Gardener OnMon The online monitoring documentation should be fixed to reflect changes to the computer configuration in the control room 2015-01-21 Complete 2015-02-23
MDR1-CMP-03 A.Dobbs OnRec The online reconstruction software should be updated to reflect recent changes to the computer configuration in the control room 2015-01-21 Updated documentation available on mice mine 2015-02-20
MDR1-CMP-04 A.Dobbs MAUS A new version of MAUS should be installed when an updated DAQ unpacking becomes available 2015-01-21 2015-03-20
MDR1-CMP-05 J.Martyniak CDB The Configuration Database get_tagged_beamline read should be fixed 2015-01-21
MDR1-CMP-06 A.Dobbs OnRec Ensure that the online reconstruction is correctly passing online reconstruction plots to the datamover directories 2015-01-21 Images being stored in standard directory on onrec02 #1639 2015-03-10
MDR1-CMP-07 P.Smith Online The web whiteboard should be mounted 2015-01-21 Complete
MDR1-CMP-08 H.Nebrensky GRID Ensure correct permissions to run reconstruction at RAL Tier1 2015-01-21 ARGUS issue reported solved 2015-02-20
MDR2-CMP-04 R.Gardener OnMon Output ROOT files of the online monitoring have to be copied at the end of the run to the agreed place. (onrec01:/home/mice/MAUS/datamover_target/) 2015-03-05 #1639
MDR2-CMP-05 A.Dobbs OnRec Output ROOT files of the online reconstruction have to be copied at the end of the run to the agreed place. (onrec01:/home/mice/MAUS/datamover_target/) 2015-03-05 #1639 2015-03-10
MDR2-CMP-06 H.Nebrensky GRID Fixed the move files script so that it picks up everything from one place (miceraid5:/data/mice/). Coordinate with A.Dobbs and R.Gardener. 2015-03-05 Complete, #1639 2015-03-10
Operations
Issue Responsible Party System Issue Description Data Noted Status Date Completed
MDR1-OPS-01 S.Boyd Documentation Check that run check lists are up to date and consistent with wiki 2015-01-21 Complete 2015-03-04